### v1

# VergeIO: Depth-Aware Eye Interaction on Glasses

Abstract.
There is growing industry interest in creating unobtrusive designs for electrooculography (EOG) sensing of eye gestures on glasses (e.g. JINS MEME and Apple eyewear). We present VergeIO, the first EOG-based glasses that enables depth-aware eye interaction using vergence with an optimized electrode layout and novel smart glass prototype. It can distinguish between four and six depth-based eye gestures with 83–98% accuracy using personalized models in a user study across 11 users and 1,320 gesture instances. It generalizes to unseen users with an accuracy of 80–98% without any calibration. To reduce false detections, we incorporate a motion artifact detection pipeline and a preamble-based activation scheme. The system uses dry sensors without any adhesives or gel, and operates in real time with 3 mW power consumption by the sensing front-end, making it suitable for always-on sensing.
copyright: acmlicensedjournalyear: 2026submissionid: 13doi: XXXXXXX.XXXXXXXconference: The 24th ACM/IEEE Conference on Embedded Artificial Intelligence and Sensing Systems; May 6–9, 2026; Irvine, USAbooktitle: The 24th ACM/IEEE Conference on Embedded Artificial Intelligence and Sensing Systems (SenSys ’26), May 6–9, 2026, Irvine, USA
1. Introduction
Figure 1. VergeIO senses eye vergence on glasses using electrooculography (EOG) by detecting gaze shifts between objects at different depths. The eyes diverge when shifting focus from a near to a far object, and converge when focusing from a far to a near object.
Given their proximity to the eyes, glasses are uniquely positioned to leverage eye movements for interaction. A less-explored movement is vergence: the inward or outward motion of the eyes that occurs naturally as humans shift their gaze between objects at different depths (Fig. 1[1]. Unlike fast, reflexive saccades or lateral gaze shifts, vergence is slower and more deliberate, making it well-suited for intentional, hands-free interaction [2, 3, 4].
Figure 2. Applications of VergeIO. (a) Varifocal lenses adjust focal depth as the user shifts gaze between ambient reference points: a nearby phone and a distant TV. (b) Smart glasses enable gaze-based selection between virtual objects at different depths. (c) Remote screening of ocular disorders using approximate reference points: thumbs and a distant object. (d) Safety monitoring for drivers by detecting abrupt vergence shifts from the road to a nearby phone.
Glasses that can sense and respond to these depth-related eye movements can enable new forms of hands-free interaction across key depth cues (Fig. 2):
  • Ambient depth cues. Varifocal glasses can dynamically refocus between a nearby book and a distant TV. A safety monitoring app on glasses can detect abrupt gaze depth shifts from the road to a phone for driver safety.
  • Virtual depth cues. Augmented reality glasses can leverage vergence to select or focus on different virtual windows.
  • Approximate physical depth cues. Remote screening for ocular disorders (e.g., convergence insufficiency, strabismus [5]) can be performed using one’s thumbs and a distant object as approximate depth markers.
Electrooculography (EOG) is a promising modality for vergence sensing on glasses, with growing industry interest in integrating EOG electrodes into glasses. Notably, these sensors have already been integrated into the nose pads and bridge of commercial-grade glasses including JINS MEME [6], and more recently, in late April 2025, Apple was granted a patent [7] to embed EOG sensors into glasses.
Unlike cameras, EOG is robust to varying ambient lighting conditions as glasses are worn across a range of environments. EOG is also low-power compared to commercial-grade cameras. In contrast to mmWave [8] and acoustic [9] systems that place sensors in front of the lenses and can obstruct the user’s vision, EOG electrodes have been integrated unobtrusively into the frames of the glasses.
The challenge, however, is that current EOG systems for glasses are not designed to robustly detect vergence. Specifically the electrodes on JINS MEME are placed too closely together to robustly detect vergence, which produces opposing electrical polarities from each eye and can cancel out or become indistinguishable [10]. As a result, such systems are typically limited to detecting blinks and coarse gaze shifts (e.g., left, right, up, down).
We present VergeIO, the first EOG-based glasses that expands the range of detectable eye movements to include distinct vergence gestures without compromising usability. Our key insight is that because EOG signals measure the eye’s electric field potential [11], reliably sensing the asymmetric movements of vergence requires a wider electrode separation. We leverage a two-channel electrode placement strategy that maximizes spatial separation, and present a novel hardware design that supports these placements, while remaining compatible with glasses.
Using this hardware setup, we show that our system can classify six vergence gestures that produce distinguishable EOG signals. These gestures correspond to three vergence distances that span everyday viewing distances: reading distance (30 cm), social interaction distance (70 cm), and far-field visual attention (200 cm).
Importantly, vergence-based interaction can in principle extend across the spectrum of glasses, from high-end augmented reality smart glasses to basic audio-only frames like Amazon Echo Frames [12], or even instrumented varifocal glasses.
For display-enabled glasses, our system can render distance markers using on-screen visual overlays. Although these markers are displayed on a fixed physical screen, they simulate depth using binocular disparity, which causes users to exhibit vergence when changing fixation between different virtual depths.
For glasses without displays, our system supports a simple physical fallback: users can use their outstretched thumbs as approximate depth reference points. One thumb is positioned at full arm’s length ($\approx$70 cm), and the other rests at half arm’s length ($\approx$30 cm), such as on the opposite elbow, along with any arbitrarily chosen far-field target. We note that the exact distance of the far-field target is less critical, as changes in vergence angle and the corresponding EOG signal are less sensitive at greater distances (Fig. 3).
Our contributions are as follows:
  • We present a depth-aware glasses system which can distinguish between four and six unique vergence gestures with an accuracy ranging from 83 to 98% in an IRB-approved study across $n=11$ users using a personalized model. Our model generalizes to unseen users with an accuracy of 80 to 98% without additional calibration.
  • A novel smart glasses prototype to sense vergence with flexible electrode contacts to accommodate different users. The prototype has a power consumption of 3 mW and can run in real-time on a lightweight microcontroller. The total weight of our system is 75 g.
  • To ensure robust performance in real-world scenarios, we design a motion artifact detection pipeline to discard unwanted events caused by facial and body movements. When evaluated over a dataset collected across 12 motion and noise classes it achieved an accuracy of 98.6% and a false negative rate of 0%, where no vergence was misclassified as noise.
  • We design a preamble-based activation scheme, where a brief eyebrow raise is performed to activate or deactivate vergence sensing. This reduces the false positive rate across several activities from 3.1% to 0%, and ensures that depth-based interactions are only initiated when intended.
  • We will open-source our hardware design and software code upon publication for the community to build on.
We envision that the sensing capability introduced by our approach can broaden the scope of eye interaction available on EOG-based glasses, enabling applications in consumer, health, and safety-monitoring domains.
2. Background and related work
Reference Sensing Modality Vergence Sensing Robust to Ambient Light Unobstructed View Glasses Form Factor Power Consumption
CIDER [13] Cameras 40 µW
Tobii [14] Cameras DNS
FocusFlow [4] Cameras DNS
Li et al. [15] Photodiodes 395 µW
GazeTrak [9] Acoustic 16.4 mW
mmET [8] mmWave 6 mW
JINS MEME [6] EOG DNS
Mishra et al. [5] EOG DNS
VergeIO (ours) EOG 3 mW
Table 1. Comparison of existing eye sensing systems on glasses. VergeIO is explicitly designed for vergence sensing using an EOG electrode configuration for glasses. (DNS = did not specify)
Eye sensing on headworn devices. There has been increased interest in eye movement sensing on head-mounted devices. Camera-based approaches [13, 16, 17, 14, 18, 19], while accurate, are power-hungry and are typically ineffective in the dark without infrared illumination. Photodiode-based systems [15], though lower in power, can be saturated by ambient sunlight in outdoor settings. Emerging acoustic [9] and mmWave systems [8] offer a promising low-power alternative, but can obstruct the user’s line of sight.
Electrode-based approaches have made their way into VR headsets [20], smart glasses [21, 22, 23, 24, 25, 26, 27, 28], and earable devices [29, 30, 31]. They are often able to concurrently record electroencephalography (EEG), electrooculography (EOG), and electromyography (EMG) signals.
Dry electrodes have been integrated into several commercial smart‐glasses platforms such as JINS MEME [6] and AttentivU [22] due to their small size, weight, and low power-consumption. They have achieved reliable blink detection and horizontal/vertical saccade tracking [10], facial‐expression recognition [24], midair gesture and context recognition [32], fatigue estimation [33], and mediated attention monitoring with real‐time biofeedback [34]. We build on this rich body of prior work by enabling vergence and depth-awareness as a key interaction primitive for EOG-based glasses.
EOG-based systems. EOG has been used to detect various eye movements, including vergence, saccades, blinks, gaze shifts, and fixations [35, 36]. Prior systems have leveraged EOG-detected eye movements as an input modality for assistive interfaces for users with physical disabilities [37], robotic control [38], and hands-free operation of wearable computers for caregivers [39]. Beyond eye movement sensing, EOG has also been used to detect facial expressions (e.g., movements of the cheek, brow, and nose) [24] and to activity recognition such as typing, reading, eating, and talking [40].
As EOG only requires small dry electrodes, there has been growing interest from both academia and industry to incorporate them into glasses [21, 27, 5]. Notable commercial efforts include JINS MEME [6], and Apple’s recently approved patent [7] to incorporate EOG electrodes into glasses. However currently proposed electrode configurations are either not able or not focused on sensing vergence, and are limited to blinks, winks, saccades and directional gaze shifts; or require a large number of electrodes that can be uncomfortable. In contrast, VergeIO expands the set of eye movements that can be detected on glasses to include six different vergence gestures using an electrode placement and a novel hardware glasses design that is optimized to capture the horizontal movements of the eye associated with vergence.
Vergence-based systems. Vergence has emerged as a reliable eye-based input modality in extended reality environments [2, 4] due to its ability to mitigate the “Midas touch” problem, which refers to the difficulty of separating deliberate gaze commands from involuntary natural exploration as the eyes move around a scene [41]. Prior work has used eye-tracking cameras to measure vergence in VR systems for selecting and activating virtual objects [2, 4], controlling linear actuators [42], and studying visual attention in children with reading difficulties [43]. EOG-based systems for vergence sensing either rely on an electrode configuration that is incompatible with the form factor of glasses [44] or rely on customized flexible electrode designs [5] that specialized fabrication, making them more challenging to scale. In contrast, our approach focuses on detecting vergence using an EOG-based electrode configuration for glasses using conventional Ag/AgCl dry electrodes.
3. System design
Figure 3. Relationship between vergence distance and angle. (a) VergeIO measures changes in vergence angle ($\theta$) using EOG signals, which vary proportionally with these changes. (b) Our system can distinguish changes in vergence angles corresponding to three interaction distances: 30, 70, and 200 cm.
3.1. Vergence distance selection
A key system design parameter is the selection of vergence distances which is the physical distance between the eye and a fixation distance. These distances must satisfy two requirements: First, they should span the range of realistic, everyday viewing distances, and second, changes in vergence between these distances should result in discernible EOG signals that can be classified.
Since EOG measures the motor movement of the eyes and directly reflects changes in vergence angle, we choose distances such that the resulting changes in vergence angle are uniformly spaced. To illustrate this, Fig. 3a shows how vergence angle ($\theta_{1}$, $\theta_{2}$, $\theta_{3}$) is defined as the angle between the lines of sight of both eyes as they converge on a fixation point.
Using an interpupillary distance (IPD) of 50 mm which is within the typical adult IPD range of 50-75 mm [45, 46], we select three vergence angles: 9.5°, 4.1°, and 1.4°. These yield uniformly spaced pairwise differences: 2.7°, 5.4°, and 8.1°, which enable them to be distinguishable in practice.
We show in Fig. 4 that the mean EOG curves for each of the six gestures when performed by using our proposed electrode placement by one participant. The figure shows that the curves are visually distinct. Additionally, we show in Fig. 3b that because the relationship between vergence angle and distance is non-linear, these vergence angles correspond to three vergence distances at key interaction distances: near (30 cm), mid (70 cm), and far (200 cm).
Figure 4. Mean horizontal EOG signal for each of the six vergence gestures for the left and right eye. Shaded region represents one standard deviation from the mean.
3.2. Placement of EOG electrodes
Electrode placement is a critical consideration in our design as it directly impacts the types of eye movements that can be detected. For instance, JINS MEME [6] is a commercial smart glass that integrates EOG sensing, but uses an electrode placement which cannot robustly sense vergence, as explicitly confirmed by prior work [10] authored by researchers from JINS MEME.
Specifically, JINS MEME places both positive electrodes on the nose pads and the negative electrode on the nose bridge with minimal spatial separation (Fig. 5a). This configuration primarily captures binocular common-mode changes rather than the asymmetric signals required to detect vergence. As vergence induces small, opposing changes in the electrical field on either side of the face, spatial separation between electrodes is critical for sensing it effectively.
To address these limitations, our system is designed to support vergence sensing while being compatible with a glasses form factor. Our design (Fig. 5b) places positive electrodes at each temple (left and right channels), a shared negative electrode on the nose bridge, and a ground electrode at the mastoid following best practices from prior work on EOG sensing [47]. By placing electrodes near the temples, our system expands the sensing area and is more sensitive to the asymmetric eye changes during vergence.
Figure 5. Limitations of prior EOG electrode placements on glasses. (a) JINS MEME’s design [6] lacks spatial separation which hinders vergence detection. Source: https://www.ohmyglasses.jp/blog/2016/05/31/jins-meme-new-world/ (b) VergeIO places electrodes at the temples and nose bridge which increases sensitivity of vergence detection. Flexible contact arms for the electrodes ensure stable skin contact across users while being compatible with a glasses form factor.
To empirically validate the advantage of our approach, we conduct a pilot study to measure the signal-to-noise ratio (SNR) of vergence gestures captured using the JINS MEME and VergeIO configuration. In this study, a single participant simultaneously wore electrodes placed according to both setups: specifically by wearing two electrodes on the temples, two on the nose pads, a shared electrode on the nose-bridge as the reference, and a ground electrode at the mastoid.
Figure 6. Hardware design. (a) Exploded CAD design illustrating the different components of the hardware design. (b) Fabricated design annotated with location of different components. (c) Front-view of glasses with dimensions.
Figure 7. SNR comparison across six vergence gestures, showing that VergeIO outperforms the commercial JINS MEME glasses [6].
We recorded four-channel EOG data over ten rounds, where each round consisted of six vergence gestures between the three distances we selected in Sec. 3.1. Synchronized data were sampled at 500 Hz and processed in sequence with a 3-order zero-phase low-pass filter at 10 Hz, a notch filter, and a third-order Savitzky–Golay filter (window size is 0.5 second). We manually segmented vergence events from the filtered signals of our two channels; for JINS MEME’s channels—where the gestures were nearly imperceptible—we used our channels’ onset–offset times to extract the corresponding segments. For each event, we extracted the two 0.2-second signals immediately before and after the event as noise segments. SNR was then computed as the ratio of the root-mean-square amplitude of the event segment to that of the concatenated noise segments.
As shown in Fig. 7, the average SNR across all six vergence gestures was 8.9 dB with our system, compared to only 1.6 dB with the JINS MEME configuration. This improvement of 7.3 dB demonstrates our system’s ability to measure vergence.
We note that while prior work [48] on EOG sensing incorporates vertical electrodes above and below each eye, vergence is primarily a horizontal movement, and vertical channels only provide modest performance gains. Additionally, we report in our evaluation that (a) the addition of vertical electrodes were reported to induce discomfort by the participants (Sec. 5.1) and (b) our electrode configuration is still able to classify between different vertical eye movements (Sec. 5.4).
3.3. Hardware design
We create a smart glasses hardware prototype to ensure reliable skin contact with our chosen electrode placement (Fig. 6):
  • Electrodes. We use dry electrodes for practicality, similar to other glasses with biosignal electrodes [6, 22]. We selected SoftPulse [49] which has been used for EOG aquisition [49, 21, 50], and is biocompatible with the skin (ISO 10993-5/10).
    • Positive electrodes. These are coupled to flexible contact arms at the user’s temple for good skin contact. These two arms can slide along the glasses’ temples to accommodate different users.
    • Reference electrode. This is mounted at the nose bridge to ensure direct contact with the user’s skin.
    • Ground electrode. This is part of a curved design at the end of the arm to maintains good contact with the user’s mastoid.
  • Amplifiers. We chose the BioAmp-EXG-Pill [51] for its compact analog front end (2.5 $\times$ 1 $\times$ 0.1 cm) with a 1000× signal gain, and its ability to interface with any 3.3 and 5 V MCU via ADC.
  • Microcontroller. We selected the Teensy 4.1 [52], integrated 12-bit ADC and high-performance Cortex-M7 core and enable precise, high-resolution bio-signal digitization and real-time processing.
  • Battery. We used the same 570 mAh rechargeable lithium battery as in Google Glass [53]. As discussed in Section  5.9, this battery enables continuous sensing for approximately 197 hours (8.2 days). We also integrated the Adafruit Micro Lipo Charger 4410 [54] into the glasses to enable on-board charging of the sensing system.
Cost. The off-the-shelf cost is \$140, with the optimized cost \$73. The cost of the ELDRY electrode coating ($4\,\mathrm{cm}^{2}$, 0.5 mm thickness) is \$20; Li-Po rechargeable battery is \$15; Li-Po micro charger is \$6; microcontroller development board is \$30; and two BioAmp-EXG-Pill amplifier modules is \$69, which can be reduced to \$2 when using the discrete components.
3.4. Motion artifact and noise removal
Figure 8. VergeIO processing pipeline.
Category Description #Segments #Recordings
Vergence Vergence 102 17
Motion Walking 30 5
Motion Turning 30 5
Motion Head tilt 30 5
Motion Stand/Sit 78 13
Motion Nodding 28 5
Noise Brow raising 48 8
Noise Saccades 62 7
Noise Blinking 42 7
Noise Chewing 42 7
Noise Talking 30 5
Mixed Chewing + Vergence 30 5
Mixed Talking + Vergence 30 5
Table 2. Composition of dataset used to train motion artifact and noise removal pipeline
As EOG signals are sensitive to facial and body movements they can obscure the vergence signal [24], it is important to identify and remove these movements to avoid false positive that classify motion or noise as vergence. These movements invoke EMG (muscular) responses that can overlap in frequency with the EOG signal, resulting in signal contamination. Because of this, these signals are also not easily separable in the frequency domain.
For this reason, we leverage a data-driven, temporal segmentation approach to identify and remove segments affected by user movement. Our noise removal pipeline is designed to operate on streaming EOG data, and operate on a moving window of 2 s with an overlap of 0.1 s. We chose 2 s as it is ample time to perform a vergence movement, which typically ranges from less than 1 s [55].
We collected a dataset (Table 2) comprising different motion and noise signals on a single participant (Fig. 9). The participant performed (1) vergence movements; (2) large body and head motions (e.g., walking, turning, head tilt, standing/sitting, nodding); (3) smaller facial and eye movements typically considered noise (e.g., brow raising, saccades, blinking, chewing, talking); and (4) noise events in combination with vergence.
For discrete movements (e.g., head tilt, stand/sit, nodding, brow raising, blinking), the participant was cued with an audible beep every 3 seconds to perform the movement. For continuous movements (e.g., walking, turning, saccades, chewing, talking), the participant sustained the activity for the full duration of each recording.
Figure 9. EOG signals recorded from the left and right eye during vergence, motion and noise events. (a) Six vergence gestures. (b) Vergence is distinct from eye movements like saccades and blinking. (c) Facial movements produce higher amplitude signals than vergence alone. (d) When combined with facial movements, vergence is dominated by facial activity, and treated as noise.
Figure 10. Experiment timeline. Participants first underwent a setup phase to ensure the glasses were properly mounted and received training on performing vergence movements along the Brock String. During the experimental phase, they completed 10 rounds, each consisting of the six vergence gestures. After completing the rounds, participants removed and remounted the glasses before repeating the procedure.
Signal conditioning. All signals were preprocessed using a third-order low-pass Butterworth filter with a 10 Hz cutoff, which was chosen conservatively to attenuate high-frequency noise while preserving the dominant spectral content of the EOG signal. A notch filter with 60 Hz was also adopted to suppress powerline interference.
Noise detection. To filter low-amplitude signals, particularly static noise from our data, we apply a dynamic amplitude threshold on our data and only process windows where samples exceed the threshold. Specifically, the threshold is computed by applying a median absolute deviation (MAD) filter over the first 2 seconds of each channel, which are assumed to be static, and multiplying this by a heuristic value of 4.5.
Signal classification. For each 2-second signal recording we computed statistical features for both left and right EOG channels, including signal mean, extrema, band power, wavelet energy, variance, RMS, peak-to-RMS ratio, trapezoidal integral, maximum derivative, and minimum derivative. We trained a logistic regression classifier to distinguish between vergence and non-vergence (motion/noise) segments. We evaluate the performance of our artifact removal pipeline in Sec. 5.7.
3.5. Vergence classification pipeline
Algorithm 1 Vergence Gesture Classifier
1:procedure ClassifyVergence($eog$)
2:     $L\leftarrow[]$
3:     for all $eg\in\text{Segment}(eog,2\text{s})$ s ) do
4:         if not IsArtifact($seg$then
5:              $seg\leftarrow\text{SavitzkyGolayFilter}(seg,250)$
6:              for all $p\in\text{DetectPeaks}(seg,30\text{mV})$ do
7:                  $g\leftarrow seg[p\!-\!0.5,p\!+\!0.5]$
8:                  $f\leftarrow\text{ExtractFeatures}(g)$
9:                  $l\leftarrow\text{RandomForest.Predict}(\text{Z-Score Normalize}(f))$
10:                  append $l$ to $L$
11:              end for
12:         end if
13:     end for
14:     return $L$
15:end procedure
16:function ExtractFeatures($g$)
17:     Split $g$ into left and right halves; for each half compute
18:      $[\max-\min,\;\int g,\;(g_{e}-g_{s})/0.5,\;\overline{\Delta g},\;\mathrm{Var}(\Delta g)]$
19:     return concatenated 10‐dim vector
20:end function
Automatic segmentation. After noise‐artifact removal, each 2 s signal window was smoothed using a Savitzky–Golay filter with a window length of 250 samples (0.5 s), which can clean the signal while preserving the features. To precisely determine the onset and offset of vergence gestures in each 2 s window, we employed the peak-detection algorithm in detecta library [56] with 30 mV as the minimum amplitude threshold, which we determined empirically from the pilot study in Sec. 3.2. Since the vergence typically range less than 1s [49], we then extracted a window centered on the detected peak, extending 0.5-seconds before and after the peak, yielding a 1 second segment.
Feature extraction. Each 1 s segment can be regarded as two halves split by the peak we detected. From each half, we extracted five features: amplitude range, definite integral, end-to-end slope, mean of the first derivative, and variance of the first derivative. This results in ten features per EOG channel. Each gestures is represented by a 20-dimensional feature vector, comprising ten features extracted from each of the two channels.
Vergence classification. Prior to model training, all features were standardized via z-score normalization to compensate for inter-user variability. The normalization also mitigates data drift due to shifts in electrode placement before and after eyeglass remounting. Using the PyCaret library [57], we automated model selection and determined that the random forest was the most effective algorithm for classifying vergence gestures.
4. Study design
4.1. Experimental setup
Figure 11. Experimental setup for user study. The stars denote the location of colored beads tied to the Brock String and used as distance markers for the vergence gestures.
Figure 12. (a) Classification accuracy ($n=11$) for within-session (training and testing on data collected without remounting the glasses) versus cross-session (training on one session and testing on the unseen session after re-mounting glasses). P10 and P11 withdrew before session 2 due to time constraints. (b) Within-session 5-fold cross-validation accuracy per participant ($n=11$) for 2-, 3- and 4-channel configurations; P10 and P11 did not complete the 3- and 4-channel trials due to discomfort.
This study was approved by our Institutional Review Board. We recruited 11 participants (aged 23.36 ± 2.06), all were free of any ocular-related disorders, including convergence insufficiency.
All the experiments were conducted in a meeting room (Fig. 11). To trigger the six vergence gestures, we used a Brock String [58], a clinical tool commonly used to assess binocular vision and diagnose convergence insufficiency. Our string includes three colored beads positioned at the designated vergence distances: red at 30 cm, green at 70 cm, and yellow at 200 cm.
In our setup (Fig. 11), participants sat in front of a laptop. The distal end of the Brock String was secured to a fixed anchor on the table at 27 cm above its surface, while the proximal end passes through a metal ring that the user holds taut at the bridge of their nose, positioned approximately 33 cm high to align with the participant’s line of sight. The string is tensioned such that it rests against the top edge of the laptop screen, helping to stabilize it throughout the task. Elevating the string in this way prevented any visual overlap of the beads, ensuring that each target remained clearly distinguishable.
4.2. Data collection protocol
We began by fitting each participant with our glasses prototype. The flexible contact arm was adjusted to ensure secure, comfortable electrode contact with the skin over the temples. We positioned the Brock String so that participants could maintain a forward gaze aligned with all three beads.
Following the standard Brock String exercise procedure [59], participants were instructed to perform each vergence gesture such that the target bead remained in sharp focus while the non-target beads appeared blurred. Our experimental protocol (Fig. 10) consisted of two sessions, each comprising ten rounds. In each round, participants performed all six vergence gestures once, cued by an audible beep, with a three-second window to complete each movement. After ten rounds, participants removed the glasses for a ten-minute break before remounting the device. For each participant, we disinfected all recording apparatus in accordance with institutional hygiene protocols.
Figure 13. (a) Confusion matrix of within-session 6 gestures classification across all the participants and sessions. (b) Confusion matrix of cross-session 6 gestures classification. (c) Confusion matrix of within-session 4 gestures classification.
5. Evaluation
5.1. Vergence classification
Within-Session Performance. We performed a within-session evaluation for each of the 11 participants in our study (Fig. 12a, Table 3), and achieve an average accuracy of 83.25 ± 5.48% across all participants at classifying between the 6 different vergence gestures. We perform a per-class analysis of these results as a confusion matrix (Fig. 13(a)). The matrix shows that near-to-far and far-to-near vergence movements are never confused, which is due to their distinctively different positive-going and negative-going shapes with opposing polarity. There is virtually no misclassification between the 30 $\rightarrow$ 200 cm and 70 $\rightarrow$ 200 cm transitions, or between the 200 $\rightarrow$ 70 cm and 200 $\rightarrow$ 30 cm transitions, which is partly due to the wider separation between their EOG curves. However, the 30 $\rightarrow$ 70 cm movement creates an EOG waveform that can overlap with the 30 $\rightarrow$ 200 cm and 70 $\rightarrow$ 200 cm movements, and the 70 $\rightarrow$ 30 cm movement likewise can overlap between the 200 $\rightarrow$ 70 cm and 200 $\rightarrow$ 30 cm movements.
We also consider an alternate version of our system using only four gestures which have a high degree of separability: (30 $\rightarrow$ 200 cm, 70 $\rightarrow$ 200 cm, 200 $\rightarrow$ 30 cm, and 200 $\rightarrow$ 70 cm). Our evaluation (Fig. 14a, Table 3) shows that the system achieves a within-session classification accuracy of 98.33%, and achieves near-perfect discrimination as shown in the confusion matrix in Fig. 13c.
Cross-Session Performance. When the user remounts the glasses, slight shifts in electrode placement introduce signal drift. To evaluate the robustness of our system against remounting, we trained a model on one of the user’s data collection sessions and tested it on the other unseen session. We then reversed the roles and averaged the results. Our evaluation (Fig.  12b, Table 3) shows that system performance reduces modestly from 83.25% to 79.91% after remounting. The confusion matrix in Fig. 13(b) shows increased misclassification between same-direction vergence movements. This is because remounting induces data drift. However, for the 4-gesture version of our system, cross-session performance remained stable at 98.33%, suggesting that signal drift from remounting was not substantial enough to impact classification accuracy.
Within Session Cross Session Cross User
6-gesture accuracy 83.25 $\pm$ 5.48% 79.91 $\pm$ 7.29% 79.72 $\pm$ 3.98%
4-gesture accuracy 98.33 $\pm$ 2.04% 98.33 $\pm$ 2.20% 98.33 $\pm$ 1.67%
Table 3. Classification performance (mean $\pm$ stdev) for 6- and 4-gesture classifier.
Cross-User Performance. Here, we evaluate the performance of a pretrained model in a leave-one-user-out setup across the 11 participants in our study, where no additional user enrollment and calibration is done. Our results in Table 3 show that in this scenario, our system achieves an accuracy of 79.72% and 98.33% without any new-user enrollment for the 6-class and 4-class version of our system respectively. Additionally, we show in Fig. 14a that these accuracies surpass the personalized model trained on 3.5 minutes of new-user data, which suggests that our system can be rapidly deployed to new users without a significant decrease in system performance.
Effect of Vertical Channels. We assessed the effect of vertically placed electrodeso on system performance. Fig. 12b shows the per-participant accuracies for within-session evaluation for our proposed two‐channel configuration, a three‐ and four-channel configuration with an additional electrode pair above and below the left and right eye. 5-fold cross-validation shows that our two-channel configuration achieved 83.25% accuracy across six gestures, while the three- and four-channel configurations result in modest performance gains of 3.09% and 3.83% and accuracies of 86.57% and 87.31% respectively. These modest gains are likely because vergence is primarily, though not exclusively, a horizontal eye movement. However, all participants reported discomfort with the 3- and 4-channel configurations during data collection, two experienced ocular discomfort discontinued data collection. These findings further support our decision to exclude vertical channels from the final design.
Figure 14. (a) Effect of training duration on classification accuracy. (b) Subgroup analysis on within-session classification accuracy. (c) Performance of the motion artifact removal pipeline on EOG segments when the user is performing vergence, motion, and noise movements. (d) False positive rate under different movement scenarios: static, walking, chewing, and talking. A preamble consisting of an eyebrow raise to activate the system reduces the false positives to zero in all four conditions.
5.2. Subgroup analysis
We performed a subgroup analysis on the within-session data collected from 11 participants across 2 sessions to evaluate system performance across different demographic groups as shown in Fig. 14(b).
Sex. When comparing the classification performance by sex, female participants ($n=3$) achieved a mean accuracy of 81.67 $\pm$ 4.38%, while male participants ($n=8$) achieved 83.22 $\pm$ 6.86%. Statistical analysis using a Student’s t-test indicated that the difference was not statistically significant (t = 0.36, p = 0.72; $\alpha$ = 0.05).
Glasses. Participants who normally wore glasses were instructed to wear them with our VergeIO glasses on top of them during the study to ensure they could clearly focus on the vergence targets at 30, 70, and 200 cm. Among participants with glasses ($n=8$) and without glasses ($n=3$), the mean classification accuracies were 83.44 $\pm$ 7.11% and 81.10 $\pm$ 1.91%, respectively. No statistically significant difference was found by Student’s t-test (t = 0.54; p = 0.60; $\alpha$ = 0.05).
5.3. Baseline comparison
We perform a comparison against two baseline EOG sensing systems.
First, we compare our system with Mishra et al.’s work [5], which is focused on vergence sensing using custom aerosol-jet–printed, stretchable, skin-like biopotential electrodes using a 2, 3, and 4-channel configurations. Their 2-channel configuration differs from ours and has a horizontal channel across both eyes, a vertical channel above and below the right eye, and a ground electrode obtrusively placed on the forehead, and reports an accuracy of 34%. In contrast, our system uses an unobtrusive 2-channel configuration using standard Ag/AgCl electrodes with the ground on the mastoid and achieves an accuracy of 83.25%. While their 3-channel and 4-channel configurations have increased accuracies of 87 and 91% respectively, additional electrodes bring about discomfort which negatively affect the user experience.
Mishra et al. [5] VergeIO (ours)
Number of electrodes 5, 5, 7 4
Number of channels 2, 3, 3 2
In-session error 34%, 87%, 91% 83.25%
Cross-session error Not specified 79.91%
Cross-user error Not specified 79.72%
Commodity electrodes
Motion artifact removal
Ground electrode Forehead (obtrusive) Mastoid (unobtrusive)
Glasses Compatibility
Angular difference tested 1–3 1.35–4.05
Table 4. Comparison between our system and the baseline proposed by Mishra et al.  [5]
Second, we compare our system’s performance against the EOG electrode layout in the commercial JINS MEME glasses [6] (Fig. 5). Here, one participant (P4) wore electrodes at the union of both our system’s and JINS MEME’s electrode positions. We then collected data across ten rounds, each including all six vergence gestures. Five‐fold cross‐validation was performed independently on both setups. Our system achieves an accuracy of 88.33 $\pm$ 1.13%, while the JINS MEME achieves a lower accuracy of 50.0 $\pm$ 1.29%. This is in line with our prior SNR analysis showing that our vergence waveforms are on average 7.3 dB higher than those from JINS MEME (Fig. 7).
5.4. Classifying vergence and non-vergence eye movements
We evaluate whether our design can classify vergence, as well as other eye movements. We collect a dataset including the four vergence gestures with high separability, along with seven additional eye movements: left and right winks, blinks, and four directional gazes (up, down, left, right). We collected 14 recordings for each class, and performed 5-fold cross validation using logistic regression. Our system achieved an accuracy of 96.1 $\pm$ 2.4%.
A few misclassifications were observed between upward and downward gaze, which is expected as our design intentionally omits vertical electrode channels to avoid user discomfort. However, our system still captures some vertical eye activity as eye movements are not perfectly directional and there are slight asymmetries in electrode placement. These results show that our electrode configuration does not tradeoff accuracy at detecting standard eye movements sensed in prior EOG systems, and can detect them robustly along with vergence gestures.
Figure 15. Confusion matrix classifying between different vergence and non-vergence movements across 11 classes.
5.5. Generalizing to virtual depth cues
So far we have demonstrated our system’s performance using real-life ambient depth cues marked by colored beads on a Brock string. Here, we evaluate whether virtual depth overlays, using stereoscopic disparity to simulate depth, can similarly induce vergence, and whether our system can accurately detect it. This setup reflects a realistic use case for smart glasses that display virtual objects at different depths.
To test, this, we developed a prototype application on the Meta Quest 2 [60], replicating the Brock string setup described in Sec.4 within a Unity-based VR environment. Beads were virtually placed at 30, 70, and 200 cm from the user’s eye center (Fig. 16a). We then invited a single participant (P4) to perform the same ten rounds of vergence gestures following the same procedure described in Sec. 4.
Using the frozen model trained on the Brock string dataset on all participants, we achieved an accuracy of 88.33% and 100% for six and four vergence-gestures classification respectively (Fig. 16c). This result shows that (1) virtual depth overlays can reliably induce vergence, even though they are rendered on a fixed display, and (2) a model trained on real-world depth cues can successfully generalize to virtual depth cues.
Figure 16. Experimental setup to evaluate generalization to (a) virtual depth cues and (b) approximate depth cues. Results are shown as confusion matrix as (c) and (d) respectively.
5.6. Generalizing to approximate depth cues
We next consider if our system can generalize to approximate physical depth cues with the user’s thumbs and a distant physical reference point (Fig. 16b). Specifically, the user fully extends one arm, rests the other hand at the elbow of the extended arm, and aligns both thumbs with a far-field object. This setup approximates the precise distances used to train the system where 30 and 70 cm, correspond approximately to half and full arm’s length for an average adult [61]. The exact depth of the far-field cue is less critical, as vergence-induced EOG signal changes diminish with increasing distance (Fig. 3).
To evaluate generalizability, a single participant (P4) performed the vergence gestures following the protocol in Sec. 4. We note that the distance between P4’s eyes and the near-, mid-, far-field markers were 20, 53, and 451 cm respectively. The minimum angular separations between these distances was 3.4° which exceeds the 2.7° separation used in the precise setup and produces separable EOG vergence waveforms.
Our cross-user model trained on data from all other participants using precise cues achieved 88.33% accuracy on the six-class task, 2.45% higher than P4’s performance with precise cues (Fig. 16d). On four class task, the model achieved 100% accuracy over the 10 rounds data. These results demonstrate that our system generalizes to scenarios without precise visual depth cues.
5.7. Motion artifact and noise removal performance
We evaluated our motion artifact and noise removal pipeline using a dataset collected from a single participant (Table 2). To avoid data leakage, we ensured that no segments from the same recording appeared in both the training and validation sets. The model achieved an average accuracy of $0.986\pm 0.009$ (Fig. 14c).
Our results show that the model correctly classified 472 of 480 noise segments, with the remaining 8 misclassified as vergence. These false positives primarily consisted of standing and saccadic movements. Below in Sec. 5.8, we demonstrate that such misclassifications can be effectively mitigated by introducing a distinct preamble to activate and deactivate the system, and thus would not pose a practical issue. Conversely, 0 of 102 vergence segments were misclassified as noise, this is because all the vergence movements exceed the amplitude threshold used to classify noise.
Figure 17. (a) Prototype of mechanically adjustable varifocal glasses with an integrated servo motor for actuation. (b) Experimental setup to determine the effective focal length at each varifocal lens setting. (c) Effective focal length of the varifocal glasses for different knob rotation ratios.
5.8. False positives from non-vergence activity
We evaluated the false positive rate of our system under four real-world conditions using a single participant: static, walking, chewing, and talking. In this context, a false positive is defined as an event incorrectly classified as a vergence gesture which can be due to shift in electrode positioning, or EMG activity that overlaps with the EOG signal frequency. We collected five recordings per condition. Static recordings lasted 5 minutes in total, during which the user fixated on a single object. Walking, chewing, and talking recordings each lasted 90 seconds. No vergence movements were performed during any of these sessions.
Using our frozen classifier trained on data described in Sec.3.4 we obtain a false positive rate across conditions rangeing from 0.43% to 5.49% (Fig. 14d). To reduce the false positive rate, we use a preamble gesture, raising the brow, as a means of activating and deactivating the system. The motion is distinct from other movements, and results in an EOG amplitude that is significantly higher than vergence (Fig. 9), and it is easy and quick to perform. After the preamble, the false positive rate reduces to 0% for all conditions. These results show that by using a brow-raise as a preamble gesture, VergeIO can significantly reduce the false positive rate.
5.9. Power consumption
We measured the power consumption of our system using the Monsoon High Voltage Power Monitor [62] at 3.7 V, over 5 runs and averaged them, with each run lasting 1 minute at the lowest clock speed of 24 MHz.
The power consumption of the two amplifier front-ends was 3.3 $\pm$ 0.004 mW. This is lower than other eye tracking systems, in particular mmWave radios [8], and microphones and speakers [9], though we note that our system is purpose-built for the specific task of vergence gesture sensing not eye-tracking.
The code for data acquisition (digitization, reading samples into buffers) consumed 4.84 $\pm$ 0.02 mW; while data processing (filtering, segmenting, classification) consumed 2.63 $\pm$ 0.02 mW.
With a 570 mAh rechargeable lithium-ion battery as in Google Glass [63], the system would operate for approximately 197 hours (8.2 days). This power draw is effectively negligible, as the overall battery life of Google Glass is dominated by other processes that resulting in a typical battery life of 8 hours [53].
We note that other standby processes on the microcontroller consume on average 138.2 $\pm$ 0.0741 mW, however we do not consider this as part of the power consumption of our proof-of-concept system, as a commercial EOG-based system would leverage a low-power ASIC or SoC, which we leave for future work.
5.10. System latency
We measured the wall-clock time of our data processing pipeline, which continuously filters and classifies the streaming EOG signal 2-second rolling windows with a 0.1-second overlap. The average processing latency across 2500 windows was 15 ± 0.002 ms, which is less than the step duration of 0.1 seconds and demonstrates our system’s ability to run in real-time.
6. Example applications
Vergence-driven varifocal glasses. We present a prototype of a mechanically tunable varifocal glass system. These glasses are useful for individuals with presbyopia, an age-related vision condition that affects more than a billion people [64] that reduces the eye’s ability to focus at different distances [65]. Importantly, while presbyopia impairs the focusing ability of the eye lens, individuals can still perform vergence movements between near and far distances [66].
Our prototype (Fig. 17a) instruments an existing pair of varifocal glasses [67] with a small servo motor [68] ($22.8\,\mathrm{mm}\times 12.2\,\mathrm{mm}\times 28.5\,\mathrm{mm}$) which is coupled to a knob to adjust the focal length. The motor weighs 29 g, which is relatively modest compared to commercial eye-tracking glasses such as the Tobii Pro Glasses 3, which weigh 388 g [14].
To show that our adjustable varifocal glasses can support focusing at distances of 30, 70, and 200 cm, we conducted a laser-based experiment to measure the effective focal lengths of the lens. The goal was to determine the focal length corresponding to different rotations of the glasses’ adjustment knob. In the setup (Fig. 17b), a laser pointer was aimed at a distant wall with the varifocal lens placed in between. Let $d$ be the distance from the laser to the lens, and $d^{\prime}$ the distance from the lens to the wall. We adjusted $d^{\prime}$ by moving the varifocal glasses until the laser spot appeared in focus, then computed the effective focal length $f$ using the lens equation [69]: $f=\frac{d\cdot d^{\prime}}{d+d^{\prime}}$. In our setup, the total distance between the laser and the wall was 12 m, which allowed us to measure focal lengths up to 3 m.
[…]
Stereoscopic depth rendering of virtual objects. Modern AR displays often render virtual objects at varying depths using stereoscopic rendering, which displays slightly different images to each eye to create a sense of depth from binocular disparity [70, 71]. Vergence sensing can be used to align virtual content with the user’s depth of focus. Importantly, although all virtual content is rendered on a screen located at a fixed physical distance, stereoscopic rendering can still induce vergence.
Figure 18. Stereoscopic depth rendering is achieved by offsetting the left and right images to simulate the three depth (30, 70, 200 cm) supported by our system.
Fig. 18 illustrates the stereoscopic depth estimation formula [72]: $D=d\cdot\left(\frac{L}{e}-1\right)$ which computes the required separation $D$ between the left and right projected images to make an object appear at a desired virtual depth $e$. Here, $L$ is the distance between the eyes and the screen, and $d$ is the interpupillary distance.
In the example, the user is positioned $L=200~{}cm$ from the screen, with an interpupillary distance of $d=5~{}cm$. To render an object at $e=30~{}cm$ and $e=70~{}cm$, the images would need to be separated by $D=28.3~{}cm$ and $D=9.3~{}cm$ respectively. While to render it at $e=200~{}cm$, which matches the physical distance to the screen, the images should overlap, resulting in zero separation.
7. Discussion
Learnability of vergence. Vergence movements have been shown to be a learnable and controllable skill even without visual reference points, enabling users to perform intentional depth-based eye gestures with training in a few minutes [73, 74, 4, 3, 75, 42]. An example voluntary vergence control is perceiving 3D autostereogram [76] using cross-eyed or wall-eyed viewing.
Eye tracking using EOG signals. While the EOG signal can estimate horizontal and vertical eye movements based on amplitude changes, it is not well suited for precise gaze tracking due to its low spatial resolution. Additionally, it captures relative changes in eye position which can lead to error accumulation over time.
However, we note that eye tracking is not an essential feature for commercial smart glasses, with devices like the Amazon Echo Frames [12] relying solely on speakers and microphones for interaction, to avoid the added cost and power demands of eye-tracking cameras.
Vergence duration. Vergence is a naturally brief process and does not require prolonged fixation by users on the close or far object. In our user study, across all participants, the EOG signal from vergence had a duration of 0.737s ± 0.170 s, without requiring extended dwell periods. This distribution suggests that vergence can enable low-latency and responsive interaction across different users.